# From Efficiency Gains to Rebound Effects: The Problem of Jevons’ Paradox in AI’s Polarized Environmental Debate

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Yet a critical dimension of AI’s climate footprint lies outside these direct resource and emissions calculations. Recent work on indirect impacts (Kaack et al., 2022) warns of potential “rebound effects”, whereby gains in efficiency spur higher overall consumption. These second-order effects challenge the presumption that purely technical optimizations alone will deliver sufficient climate benefits. Cost savings achieved by more efficient AI hardware, for example, can spur increased demand for new AI functionalities, which in turn drive further hardware upgrades. Economists refer to such transformations as Jevons’ Paradox, which was proposed in the 19th century by economist William Stanley Jevons, who observed that as coal use became more efficient, it was also paradoxically leading to an increase, and not a decrease, in the consumption of coal across different industries (Jevons, 1866). The addition of increasingly efficient AI to systems from commerce to transportation can have far-reaching effects on our societies, our behaviors, and the future paths available to us in the race against climate change.
@@ 2.1. Arguments that AI is a net climate negative
Data from the International Energy Agency (IEA) has provided the strongest international benchmark. They note that the electricity demand from data centers, driven heavily by AI training and inference, is currently at 2% of total global electricity and will more than double by 2026, surpassing Canada’s national power use (International Energy Authority, 2023; Guidi et al., 2024). This growth is putting strain on energy grids around the world, resulting in many countries, such as Ireland and the Netherlands, placing a moratorium on the construction of new data centers in certain regions. In the United States, the demand for electricity has surged in the past twelve months – according to a recent report into the national grid, electric utilities have nearly doubled their forecasts of how much additional power they will need in the next five years (Wilson and Zimmerman, 2023).
@@
AI technologies are currently extensively used in oil and gas exploration and drilling, significantly improving efficiency and increasing yield (Sircar et al., 2021; Tariq et al., 2021). These enabled emissions are difficult to quantify due to lack of corporate transparency and reporting, but they have been attracting scrutiny in recent years, given the environmental harm that they perpetuate. For instance, a coalition of Microsoft employees estimated that a single deal the company struck with Exxon Mobil that uses AI to expand oil and gas production in Texas and New Mexico by 50,000 barrels of oil per day could add up to 640 percent compared to the company’s carbon removal targets for the year (Stone, 2024), yet these numbers were not included in the company’s carbon accounting and reporting efforts (Stone, 2020). Other firms have been equally opaque about their AI-driven enabling of the oil and gas exploration, but a Greenpeace report from 2020 ties all of the major technology companies to the sector (GreenPeace, 2020).
@@
AI’s expanding operational footprint also contributes to electronic waste (e-waste), which is now the fastest-growing segment of solid waste worldwide, reaching 62 million tonnes in 2022. Ongoing reports have pointed to the urgent need to find ways to reduce and reuse this material more effectively given its status as both an environmental and health hazard (Robinson, 2009). The UN’s Global E-Waste Monitor 2024 showed that about 22% of e-waste has been shown to be formally collected and recycled, with global generation of electronic waste rising five times faster than e-waste recycling (Baldé et al., 2024). The remainder ends up dumped in landfills, often in developing countries, where researchers assess both how hazardous substances like mercury, arsenic, and lead leach into local ecosystems and how they impact public health. High turnover in AI hardware is accelerating e-waste output: Although GPUs can last about five years, the push for higher performance is prompting more frequent upgrades – one recent study estimates that AI will generate an additional 1.2–5 million metric tons of e-waste by 2030 (Wang et al., 2024).
@@ 2.2. Arguments that AI is a net climate positive
The positive potential of AI for climate action has generated considerable enthusiasm, and there are many existing and in-progress projects and initiatives that aim to harness this potential into concrete tools and applications (Rolnick et al., 2022; Vinuesa et al., 2020). There are several key areas in which this is particularly salient, from generating insights from large quantities of data from different modalities to optimizing complex systems and tools (Clutton-Brock et al., 2021). Concretely, this can be seen in applications of AI that analyze satellite imagery to identify deforestation (Vorotyntsev et al., 2021), methane leaks (Schuit et al., 2023), or even the health of coral reefs from space (Burns et al., 2022). AI has also improved the accuracy of weather forecasts, which can help renewable energy grid operators predict the output of solar panels and wind turbines, which can help streamline the transition towards renewable energy sources on a global scale (Sharma et al., 2011; Krechowicz et al., 2022) In recent years, the development of AI for accelerating scientific discovery in materials science has become an area of focus especially in applying generative AI models such as LLMs, which are able to assess or generate potential new compounds with climate-positive applications such as renewable energy and carbon capture (Bengio et al., 2023; Merchant et al., 2023; Zeni et al., 2025). Finally, the use of AI in solar geoengineering (i.e. injecting aerosol particles into the atmosphere to reflect sunlight and reduce global warming) has also gathered traction, given AI’s potential to help predict the impacts of geoengineering on existing ecosystems and avoid possible negative side effects (Agrawal and de Witt, 2025).
@@
AI-driven products and services are the primary contributor to AI’s direct environmental impacts, and leading AI companies including Google, Microsoft, Amazon and Meta have committed to achieving net-zero carbon emissions for their business operations, including AI development and use, within the next 5-15 years (Google, 2024; Microsoft, 2024; Meta, 2021; Amazon Web Services, 2021). The primary mechanism by which these companies plan to achieve their goal is by increasing the use of renewable energy over that period, alongside compensating for emissions using market-based mechanisms such as power purchase agreements (PPAs), carbon offsets, and investment in carbon-reducing technologies. The high-level idea behind these market-based mechanisms is that any carbon-producing activity can be negated by an equally carbon-reducing activity, resulting in net zero emissions globally. In the absence of significant advances in renewable energy production, storage and distribution (which, like other transformative changes in the climate space, would also require significant investment in infrastructure and navigating corresponding sociopolitical systems), PPAs and carbon offsets will remain necessary to fill the gap. However, offsetting was only ever meant to serve as a temporary stop-gap to help reduce emissions in the short term, and does not represent a viable replacement to reducing actual emissions. Fundamental limitations to carbon offsetting are: (1) the difficulty of proving additionality, or that the carbon-reducing activity would not have happened regardless of investment or purchase of the offset (Campbell et al., 2018; Hahn and Richards, 2013), and (2) offsetting rarely mitigates localized impacts to the communities where emissions or other environmental degradation is occurring. Increasingly, companies such as Intel and TSMC are reporting annual water withdrawal and consumption as well, and compensating with market-based water offsets which are subject to the same challenges (Woodward et al., 2016).
@@
Similarly to Jevons’ Paradox, just because an AI model becomes more efficient, that does not imply that overall AI resource consumption will decrease, and in fact the inverse effect is highly plausible. Koomey and Masanet (2021) cogently argue that this is not the first time the alarm has been raised about rising energy use due to technology, and that similar projections made in the dot-com boom of the early 2000s failed to materialize. They cite poor data availability, flawed methodology, and inaccurate reporting as causes for inaccurate projections in data center energy use that ultimately did not take into account significant improvements in data center efficiency. Similar uncertainty is clouding the prediction of AI’s energy use, and we are well aligned with Koomey and Masanet (2021), Masanet et al. (2020a) and others in calling for more granular transparent data from technology firms and service providers, and more rigorous analysis of available data. This does not imply, however, that AI will necessarily follow the same trend as past technological advances. Further, although data center energy use is a relatively accessible statistic for approximating the environmental impacts of ICT, it represents only small portion of the diverse direct and indirect negative environmental impacts arising from AI, which extend far beyond GHG emissions due to the energy required to develop and use AI models.
@@
Some argue that there is no reason to raise concerns around the current or future energy use (and corresponding environmental impacts) of AI specifically, as compared to any other way that energy might be used, such as “for watching television, microwaving popcorn, [or] powering lights” (Castro, 2024). This line of thought posits that AI will be subject to the same market pressures, such as energy prices, as any other use case, and as a result will benefit from the same innovations, such as a market transitions to renewable energy. This oversimplification ignores the reality that (1) AI is already responsible for non-trivial negative environmental externalities and (2) we can, and likely should, regulate certain uses of AI as appropriate/inappropriate or necessary/unnecessary, and that this does require breaking down the monolith of “AI” into different use cases and corresponding judgments of potential utility. For example, under a limited energy or emissions budget, we might selectively incentivize uses of AI that contribute towards the UN’s sustainable development goals (UN General Assembly, 2015) or mitigate national security concerns, over e.g. generating personalized ads for social media (as we discussion in Section 3). Finally, AI does differ substantially from other energy sinks in its potential to engender vast transformations in the economy and society, similar to how the advent of the Internet has undeniably changed the nature of work, education, and social interactions. Purely economic incentives have obviously failed to align well with environmental sustainability in the past, and AI is no different; this does not mean that we should not strive to do better with this new technology.
While the debates over AI’s role in climate change and sustainability have become increasingly polarized, both sides have tended to focus on the direct impacts of this technology with both sides concentrating on the direct impact— positive and negative. But direct effects are not the whole story. As we show in the Section below, the integration of AI into tools and systems reshapes social structures and influences human behavior, which ultimately has complex environmental consequences.
@@ 3. Indirect Impacts and Rebound Effects
Scale effects AI models growing in size and complexity Optimization approaches for AI training and deployment (batching, caching)
@@
Indirect economic rebound effects More consumer devices (speakers, appliances, etc.) that allow users to interact with AI systems Smartphones replaced more often to run newest AI functionalities
Economy-wide rebound effects Power purchase agreements from AI compute providers enable global renewable energy capacity to be expanded
@@ 3.1. Material Rebound Effects
For instance, substitution impacts arise when one product or service replaces another, rendering digital something that was previously analog (i.e. dematerializing it). Recent examples of this include online streaming which has replaced VHS cassettes, vinyl records, CDs and DVDs, as well as e-readers substituting print books and magazines (Hendrickson et al., 2010). This can come with both positive and negative impacts in terms of sustainability: Increased capacity and efficiency of new technologies (e.g. many different movies and TV shows being available on a single streaming platform) can be invalidated by the increased materials needed to support the underlying infrastructure needed to host and deliver these technologies (Rieger, 2021). For instance, an LCA has been carried out comparing print books versus e-readers, finding that 115 books would produce the same amount of CO2 as one Amazon Kindle (Dowd-Hinkle, 2012; Scrogum, 2009). In terms of AI, many previously material tools and products are being substituted by digital, AI-enabled systems. For instance, paper maps have all but been replaced by navigation apps for daily commutes and activities. In other cases, the substitution is gradual. For instance, only a fraction of illustrations such as photographs and artworks have, as of yet, been replaced by AI-generated images, and combinations of both AI tools and analog dictionaries are used for translation and writing tasks. While these substitution impacts are likely environmentally-positive (given that it is no longer necessary to manufacture the equipment and supplies needed to carry out the original task), the environmental impacts of developing and deploying these AI-enabled tools remains under-explored and direct comparisons are seldom performed.
@@
These kinds of shifts in the way objects and processes operate can, in turn, come with space rebound effects that change the way physical space is used with the introduction of a new product. For instance, new approaches such as e-commerce have manifested changes in the physical space occupied by brick-and-mortar stores and warehouses, resulting in smaller stores and bigger warehouses (Tiwari and Singh, 2011). This, in turn, has impacts on and supply chain dynamics, redistributing the environmental footprint of operations (e.g. heating and cooling), logistics and transportation  (Girod et al., 2011). Similarly, the proliferation of videoconferencing platforms and the increased accessibility of high-speed internet, which facilitate remote work, also comes with space rebound effects, with more people working from home, requiring home offices (and therefore bigger homes), and less need for dedicated office space in commercial buildings, with further consequences in urban planning, mobility, and building maintenance  (Hostettler Macias et al., 2022). Finally, the size of digital devices has been steadily shrinking over the last decade. For instance, the average mass of a mobile phone shrank by half in the last decade (Hilty et al., 2006), whereas the average size of data centers has doubled (Statista, 2023). With the advent of user-facing AI tools and services running on shrinking mobile devices, the configuration of data centers themselves has shifted from a distributed network of smaller, in-house server rooms to warehouse-sized hyperscale data centers that have tens of thousands of servers in one place. These large, monolithic concentrations of servers enable the large-scale training of AI models, since it allows the high-speed interconnection of thousands of compute nodes serving as a single supercomputer (Hoefler et al., 2022). While hyperscale data centers typically achieve the highest efficiency (measured as PUE), their resource consumption intensity, further magnified by their exceptionally high geographic concentration (International Energy Authority, 2024), can put undue strain on local infrastructure. For example, Loudoun County in the U.S. State of Virginia has become known as “Data Center Alley,” hosting 25 million square feet of data centers that use a quarter of the state’s energy and put extreme strain on its energy supply (Tremayne-Pengelly, 2024).

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournalyear: 2025copyright: ccconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732007isbn: 979-8-4007-1482-5/2025/06
@@ 1. Introduction
Yet a critical dimension of AI’s climate footprint lies outside these direct resource and emissions calculations. Recent work on indirect impacts (Kaack et al., 2022) warns of potential “rebound effects”, whereby gains in efficiency spur higher overall consumption. These second-order effects challenge the presumption that purely technical optimizations alone will deliver sufficient climate benefits. Cost savings achieved by more efficient AI hardware, for example, can spur increased demand for new AI functionalities, which in turn drive further hardware upgrades and increase costs. Economists refer to such transformations as Jevons’ Paradox, which was proposed in the 19th century by economist William Stanley Jevons, who observed that as coal use became more efficient, it was also paradoxically leading to an increase, and not a decrease, in the consumption of coal across different industries (Jevons, 1866). The addition of increasingly efficient AI to systems from commerce to transportation can have far-reaching effects on our societies, our behaviors, and the future paths available to us in the race against climate change.
@@ 2.1. Arguments that AI is a net climate negative
Data from the International Energy Agency (IEA) has provided the strongest international benchmark. They note that the electricity demand from data centers, driven heavily by AI training and inference, is currently at 2% of total global electricity and will more than double by 2026, surpassing Canada’s national power use (International Energy Authority, 2023; Guidi et al., 2024). This growth is putting strain on energy grids around the world and resulting in many countries, such as Ireland and the Netherlands, placing a moratorium on the construction of new data centers in certain regions. In the United States, the demand for electricity has surged in the past twelve months – according to a recent report into the national grid, electric utilities have nearly doubled their forecasts of how much additional power they will need in the next five years (Wilson and Zimmerman, 2023).
@@
AI technologies are currently extensively used in oil and gas exploration and drilling, significantly improving efficiency and increasing yield (Sircar et al., 2021; Tariq et al., 2021). These enabled emissions are difficult to quantify due to lack of corporate transparency and reporting, but they have been attracting scrutiny in recent years, given the environmental harm that they perpetuate. For instance, a coalition of Microsoft employees estimated that a single deal the company struck with Exxon Mobil that uses AI to expand oil and gas production in Texas and New Mexico by 50,000 barrels of oil per day could add up to 640 percent more carbon emissions compared to the company’s carbon removal targets for the year (Stone, 2024), yet these numbers were not included in the company’s carbon accounting and reporting efforts (Stone, 2020). Other firms have been equally opaque about their AI-driven enabling of the oil and gas exploration, but a Greenpeace report from 2020 ties all of the major technology companies to the sector (GreenPeace, 2020).
@@
AI’s expanding operational footprint also contributes to electronic waste (e-waste), which is now the fastest-growing segment of solid waste worldwide, reaching 62 million tonnes in 2022. Ongoing reports have pointed to the urgent need to find ways to reduce the amount of waste and to reuse these materials more effectively given the status of e-waste as both an environmental and health hazard (Robinson, 2009). The UN’s Global E-Waste Monitor 2024 showed that about 22% of e-waste has been shown to be formally collected and recycled, with global generation of electronic waste rising five times faster than e-waste recycling (Baldé et al., 2024). The remainder ends up dumped in landfills, often in developing countries, where researchers assess both how hazardous substances like mercury, arsenic, and lead leach into local ecosystems and how they impact public health. High turnover in AI hardware is accelerating e-waste output: although GPUs can theoretically last about five years, the push for higher performance is prompting more frequent upgrades – one recent study estimates that AI will generate an additional 1.2–5 million metric tons of e-waste by 2030 (Wang et al., 2024).
@@ 2.2. Arguments that AI is a net climate positive
The positive potential of AI for climate action has generated considerable enthusiasm, and there are many existing and in-progress projects and initiatives that aim to harness this potential into concrete tools and applications (Rolnick et al., 2022; Vinuesa et al., 2020). There are several key areas in which this is particularly salient, from generating insights from large quantities of data in different modalities to optimizing complex systems and tools (Clutton-Brock et al., 2021). Concretely, this can be seen in applications of AI that analyze satellite imagery to identify deforestation (Vorotyntsev et al., 2021), methane leaks (Schuit et al., 2023), or even the health of coral reefs from space (Burns et al., 2022). AI has also improved the accuracy of weather forecasts, which can help renewable energy grid operators predict the output of solar panels and wind turbines, which can help streamline the transition towards renewable energy sources on a global scale (Sharma et al., 2011; Krechowicz et al., 2022) In recent years, the development of AI for accelerating scientific discovery in materials science has become an area of focus, especially in applying generative AI models, which are able to assess or generate potential new compounds with climate-positive applications such as renewable energy and carbon capture (Bengio et al., 2023; Merchant et al., 2023; Zeni et al., 2025). Finally, the use of AI in solar geoengineering (i.e. injecting aerosol particles into the atmosphere to reflect sunlight and reduce global warming) has also gathered traction, given AI’s potential to help predict the impacts of geoengineering on existing ecosystems and avoid possible negative side effects (Agrawal and de Witt, 2025).
@@
AI-driven products and services are the primary contributor to AI’s direct environmental impacts, and leading AI companies including Google, Microsoft, Amazon and Meta have committed to achieving net-zero carbon emissions for their business operations, including AI development and use, within the next 5-15 years (Google, 2024; Microsoft, 2024; Meta, 2021; Amazon Web Services, 2021). The primary mechanism by which these companies plan to achieve their goal is by increasing the use of renewable energy over that period, alongside compensating for emissions using market-based mechanisms such as power purchase agreements (PPAs), , long-term contracts between energy providers and customers who promise to purchase a certain amount of generated energy over a decades-long horizon, carbon offsets, and investment in carbon-reducing technologies. The high-level idea behind these market-based mechanisms is that any carbon-producing activity can be negated by an equally carbon-reducing activity, resulting in net zero emissions globally. In the absence of significant advances in renewable energy production, storage and distribution (which, like other transformative changes in the climate space, would also require significant investment in infrastructure and navigating corresponding sociopolitical systems), PPAs and carbon offsets will remain necessary to fill the gap. However, offsetting was only ever meant to serve as a temporary stop-gap to help reduce emissions in the short term, and does not represent a viable replacement to reducing actual emissions. Fundamental limitations to carbon offsetting are: (1) the difficulty of proving additionality, or that the carbon-reducing activity would not have happened regardless of investment or purchase of the offset (Campbell et al., 2018; Hahn and Richards, 2013), and (2) that offsetting rarely mitigates localized impacts to the communities where emissions or other environmental degradation is occurring. Increasingly, companies such as Intel and TSMC are reporting annual water withdrawal and consumption as well, and compensating with market-based water offsets which are subject to the same challenges (Woodward et al., 2016).
@@
Similarly to Jevons’ Paradox, just because an AI model becomes more efficient, that does not imply that overall AI resource consumption will decrease, and in fact the inverse effect is highly plausible. However, as Koomey and Masanet (2021) cogently argue, this is not the first time the alarm has been raised about rising energy use due to technology, and that similar projections made in the dot-com boom of the early 2000s failed to materialize. They cite poor data availability, flawed methodology, and inaccurate reporting as causes for inaccurate projections in data center energy use that ultimately did not take into account significant improvements in data center efficiency. Similar uncertainty is clouding the prediction of AI’s energy use, and we are well aligned with Koomey and Masanet (2021), Masanet et al. (2020a) and others in calling for more granular transparent data from technology firms and service providers, and more rigorous analysis of available data. This does not imply, however, that AI will necessarily follow the same trend as past technological advances. Further, although data center energy use is a relatively accessible statistic for approximating the environmental impacts of ICT, it represents only small portion of the diverse direct and indirect negative environmental impacts arising from AI, which extend far beyond GHG emissions due to the energy required to develop and use AI models.
@@
Some would argue that there is no reason to raise concerns around the current or future energy use (and corresponding environmental impacts) of AI specifically, as compared to any other way that energy might be used, such as “for watching television, microwaving popcorn, [or] powering lights” (Castro, 2024). This line of thought posits that AI will be subject to the same market pressures, such as energy prices, as any other use case, and as a result will benefit from the same innovations, such as a market transitions to renewable energy. This oversimplification ignores the reality that (1) AI is already responsible for non-trivial negative environmental externalities and (2) we can, and likely should, regulate certain uses of AI as appropriate/inappropriate or necessary/unnecessary, and that this does require breaking down the monolith of “AI” into different use cases and corresponding judgments of potential utility. For example, under a limited energy or emissions budget, we might selectively incentivize uses of AI that contribute towards the UN’s sustainable development goals (UN General Assembly, 2015) or mitigate national security concerns, over e.g. generating personalized ads for social media (as we discussion in Section 3). Finally, AI does differ substantially from other energy sinks in its potential to engender vast transformations in the economy and society, similar to how the advent of the Internet has undeniably changed the nature of work, education, and social interactions. Purely economic incentives have obviously failed to align well with environmental sustainability in the past, and AI is no different; this does not mean that we should not strive to do better with this new technology.
While debates over AI’s role in climate change and sustainability have become increasingly polarized, both sides have tended to focus only on the direct impacts of this technology— positive and negative. But direct effects are not the whole story. As we show in the next section, the integration of AI into tools and systems reshapes social structures and influences human behavior, which ultimately has complex environmental consequences.
@@ 3. Indirect Impacts and Rebound Effects
Scale effects AI models growing in size and complexity
@@
Indirect economic rebound effects More consumer devices (speakers, appliances, etc.) that allow users
to interact with AI systems
Economy-wide rebound effects Power purchase agreements from AI compute providers enable global
renewable energy capacity to be expanded
@@ 3.1. Material Rebound Effects
For instance, substitution impacts arise when one product or service replaces another, rendering digital something that was previously analog (i.e. dematerializing it). Recent examples of this include online streaming which has replaced VHS cassettes, vinyl records, CDs and DVDs, as well as e-readers substituting print books and magazines (Hendrickson et al., 2010). This can come with both positive and negative impacts in terms of sustainability: Increased capacity and efficiency of new technologies (e.g. many different movies and TV shows being available on a single streaming platform) can be invalidated by the increased materials needed to support the underlying infrastructure needed to host and deliver these technologies (Rieger, 2021). For instance, a life cycle assessment (LCA), which evaluates the environmental impacts of an artifact arising throughout its existence (typically including disposal), has been performed comparing print books to e-readers, finding that 115 books would produce the same amount of CO2 as a single Amazon Kindle device (Dowd-Hinkle, 2012; Scrogum, 2009). In the case of AI, many previously material tools and products are being substituted by digital, AI-enabled systems, such as paper maps having been replaced by digital navigation for routine travel. Substitution impacts may occur at varying speeds and may be more gradual. For instance, only a fraction of illustrations such as photographs and artworks have, as of yet, been replaced by AI-generated images, and combinations of AI tools and analog dictionaries are used for translation and writing tasks. While these substitution impacts are likely environmentally-positive (given that it is no longer necessary to manufacture the equipment and supplies needed to carry out the original task), the environmental impacts of developing and deploying these AI-enabled tools remains under-explored and direct comparisons are seldom performed, or even possible using status quo data and methodology (see Section 3.4).
@@
These kinds of shifts in the way objects and processes operate can, in turn, come with space rebound effects that change the way physical space is used with the introduction of a new product. For instance, new approaches such as e-commerce have manifested changes in the physical space occupied by brick-and-mortar stores and warehouses, resulting in smaller stores and bigger warehouses (Tiwari and Singh, 2011). This, in turn, has impacts on supply chain dynamics, redistributing the environmental footprint of operations (e.g. heating and cooling), logistics and transportation  (Girod et al., 2011). Similarly, the proliferation of videoconferencing platforms and the increased accessibility of high-speed internet, which facilitate remote work, also comes with space rebound effects, with more people working from home, requiring home offices (and therefore bigger homes), and less need for dedicated office space in commercial buildings, with further consequences in urban planning, mobility, and building maintenance  (Hostettler Macias et al., 2022). Finally, the size of digital devices has been steadily shrinking over the last decade. For instance, the average mass of a mobile phone shrank by half in the last decade (Hilty et al., 2006), whereas the average size of data centers has doubled (Statista, 2023). With the advent of user-facing AI tools and services running on shrinking mobile devices, the configuration of data centers themselves has shifted from a distributed network of smaller, in-house server rooms to warehouse-sized hyperscale data centers that have tens of thousands of servers in one place. These large, monolithic concentrations of servers enable the large-scale training of AI models, since it allows the high-speed interconnection of thousands of compute nodes serving as a single supercomputer (Hoefler et al., 2022). While hyperscale data centers typically achieve the highest efficiency (measured as PUE), their resource consumption intensity, further magnified by their exceptionally high geographic concentration (International Energy Authority, 2024), can put undue strain on local infrastructure. For example, Loudoun County in the U.S. State of Virginia has become known as “Data Center Alley,” hosting 25 million square feet of data centers that use a quarter of the state’s energy and put extreme strain on its energy supply (Tremayne-Pengelly, 2024).